OpenStudy (anonymous):

numbers and integers: find three consecutive odd integers such that the sum of the first and third is equal to 27 less than 3 times the second.

OpenStudy (anonymous):

x + x + 4 = 3(x + 2) - 27 2x + 4 = 3x + 6 - 27 -x = -25 x = 25 25, 27, 29